The Gamma Corporation is financed 75% with common stock and 25% with debt. The debt currently has an interest rate, before tax, of 11% per year. Gamma’s common stock trades at $20 per share. The most recent quarterly dividend was $0.60. Future dividends are expected to grow 1.2% per quarter. Gamma’s marginal tax rate is 39%. Find the annual cost of equity. (Hint: Compute an EAR) Question 6 options: 17.55% 16.39% 19.32% 18.05%
Answer #1

Quarterly rate = Dividend *(1+growth)/Price + growth = 0.60*(1+1.2%)/20 + 1.2% = 4.236%
EAR = (1+4.236%)^4 -1 = 18.05%

So option d is correct

Please Discuss in case of Doubt

Best of Luck. God Bless
Please Rate Well